# Large Spatial Model: End-to-end Unposed Images to Semantic 3D

### v1

@@ Abstract
In this work, we introduce the Large Spatial Model (LSM), which directly processes unposed RGB images into semantic radiance fields. LSM simultaneously estimates geometry, appearance, and semantics in a single feed-forward pass and can synthesize versatile label maps by interacting through language at novel views. Built on a general Transformer-based framework, LSM integrates global geometry via pixel-aligned point maps. To improve spatial attribute regression, we adopt local context aggregation with multi-scale fusion, enhancing the accuracy of fine local details. To address the scarcity of labeled 3D semantic data and enable natural language-driven scene manipulation, we incorporate a pre-trained 2D language-based segmentation model into a 3D-consistent semantic feature field. An efficient decoder parameterizes a set of semantic anisotropic Gaussians, allowing supervised end-to-end learning. Comprehensive experiments on various tasks demonstrate that LSM unifies multiple 3D vision tasks directly from unposed images, achieving real-time semantic 3D reconstruction for the first time.
@@ 1 Introduction
Furthermore, reasoning about and interacting with the environment would benefit from a comprehensive 3D understanding. Open-vocabulary methods, which perform semantic segmentation without relying on a fixed set of labels, provide notable flexibility. However, unlike single-image understanding, the absence of large-scale and diverse 3D scene data with accurate multiview language annotations complicates the challenge. Efforts have been made to integrate 2D features into frameworks such as NeRF [4, 12, 13] and 3D-GS [5, 14, 15]. Yet, these methods, such as Feature-3DGS [14] and DreamScene360 [15], typically require overfitting each 3D scene separately with extensive captured viewpoints and preprocessing camera poses using Structure-from-Motion.
To address the challenges outlined above, we propose for the first time a novel unified framework for these key 3D vision subproblems: dense 3D reconstruction, open-vocabulary semantic segmentation, and novel view synthesis from unposed and uncalibrated images. Our approach leverages a single Transformer-based model that learns the attributes of a 3D scene via a point-based semantic radiance field. Unlike previous methods that rely on epipolar Transformers with known camera parameters [16, 17, 18] or require extensive per-scene fitting [5, 14], we employ a coarse-to-fine strategy. This strategy predicts dense 3D geometry using pixel-aligned point maps, progressively refining these points into anisotropic Gaussians in a single feed-forward pass.
@@ 2 Related Work
Structure-from-Motion (SfM) aims to jointly estimate camera poses and reconstruct sparse 3D structures from multiple views. Traditional pipelines [1] involve multiple stages, including descriptor extraction, correspondence estimation, and incremental bundle adjustment. Recent advances in learning-based techniques [20, 21, 22, 23, 24] have further improved the accuracy and efficiency of SfM. These methods are widely adopted in 3D vision tasks, where differentiable neural representations typically assume accurate camera poses provided by SfM. For instance, NeRF [4] and its successors [25, 26] rely on poses estimated offline via COLMAP [1, 27]. Similarly, 3D Gaussian Splatting [5] uses SfM-generated 3D points for initialization. Beyond novel view synthesis, lifting 2D features to 3D has gained traction in various editing tasks [12, 14, 28, 13].
@@
3D reconstruction is a long-standing problem in computer vision, with traditional approaches like SfM [29, 30, 1], Multi-view Stereo (MVS) [3, 2, 31, 32], and Signed Distance Function (SDF) [33, 34]. More recent techniques utilize neural representations, including implicit [4] and explicit [5] formats to generate 3D models. Semantic understanding is often integrated during the reconstruction process [35], or through additional optimization steps [36, 12]. However, most methods depend on a preprocessing step like SfM [1] to estimate camera calibration, poses, and sparse point clouds before dense reconstruction, either through feed-forward prediction or test-time optimization. This reliance on calibration and pose estimation limits scalability with large-scale data, contrasting the success seen with large foundation models.
The latest pose-free, feedforward approaches, such as DUSt3R [19], MASt3R [37], and InstantSplat [38], bypass these limitations by predicting dense point clouds directly from unposed stereo image pairs, allowing pixel-aligned geometry prediction at normalized scales. These methods benefit from training on diverse and massive 3D datasets.
In contrast, our framework offers a holistic solution for dense 3D semantic reconstruction from unposed images. It integrates dense 3D geometry reconstruction, novel view synthesis, and language-based 3D interaction, while minimizing the need for extensive data annotation. Since dense 3D annotations are often scarce in real-world scenarios, we leverage semantic anisotropic Gaussians to lift 2D features to 3D with minimal annotation. Our approach addresses more advanced "high-level" 3D tasks than DUSt3R [19], solving them jointly within a unified framework.
@@ 3.1 Dense Geometry Prediction
where is pixel-aligned confidence map, same as DUSt3R, indicates all valid points to the origin, is a hyper-parameter that promotes regularization, encouraging the network to perform robustly in challenging areas. The depth error is calculated by
where the normalization operation indicates that the predicted and ground-truth pointmaps are processed by scaling factors and , respectively, which simply represent the average distance.
@@ 3.2 Point-wise Feature Aggregation
Building on the foundational work in NeRF [4] and multi-view stereo [3], which employ a coarse-to-fine strategy for high-quality radiance field and depth estimation, we extend this approach by applying a Transformer at the point level. This Transformer utilizes hierarchical representations to achieve finer, point-based predictions.
@@
Rather than relying solely on a single network to represent the scene, we employ two Transformer-based networks optimized for distinct tasks: one for capturing "coarse" global geometry and another for "fine" local information aggregation. Initially, we integrate stereo point maps, including color information for each point primitive, formulated as to serve as input. Unlike tokenized image patches, point primitives carry distinct geometric significance within Euclidean space. Inspired by recent advancements in point-cloud processing [43, 44, 45], we employ a Transformer within a localized window to perform point-wise aggregation, selectively emphasizing key features from neighboring primitives. Point-wise encoding and decoding are essential for refining scene representation, utilizing multiscale aggregation across five hierarchical levels.
@@
The original point features contain explicit and precise spatial information, which is critical for accurate geometry reconstruction. In contrast, the image token features are rich in semantic content, providing important contextual information that enhances general understanding of the scene. Cross-model fusion enables the integration of detailed spatial geometry with semantic richness:
@@ 3.3 Learning Hierarchical Semantics
To facilitate semantic 3D representation, we augment the anisotropic 3D Gaussians with a learnable semantic feature embedding (a.k.a. semantic anisotropic Gaussians) and rasterize the 3D structure into the 2D image plane by blending Gaussians that overlap with each pixel using a feature rendering function. After obtaining the embeddings on 2D images, we optimize the feature vectors on the 3D Gaussian by minimizing the difference between the rasterized feature map and the feature maps generated by a pre-trained 2D model. Unlike the previous method [14] which requires test time optimization, we transform the learning of the feature field into a fully learnable process.
@@
Feature maps from a pre-trained 2D model are inherently view-inconsistent due to the lack of spatial awareness during the 2D model’s training. To elevate multi-view feature embeddings into a coherent 3D feature field for holistic 3D understanding, we introduce a dynamic fusion strategy employing an attention-based correlation module. This module is specifically designed to learn blending weights for each point-based primitive from the input pixel-wise feature embedding. We employ attention blocks as described in Eq.3.2 to synchronize the latent spaces of semantic image tokens from stereo inputs with point-based networks through a supplementary set of cross-attention layers. The visual feature from LSeg[46], denoted as , is utilized for this purpose.
This loss function is minimized during training by utilizing rasterized feature maps on new views and directly inferred feature maps using ground truth images on new views (LSeg [46]), thereby transferring knowledge and facilitating the lifting of the feature field.
@@
To improve model efficiency, we propagate information from (ViT feature) and the semantic feature from the frozen Lseg , to (point feature), which has fewer tokens, thereby enabling selective attention to critical features. We further refine feature fusion across multiple stages, optimizing information flow while minimizing additional computational overhead.
@@ 3.4 Training Objective
where and are rasterized and GT pixel intensities, denotes represented 3D scene using a set of 3D Gaussians, and denotes rendered LSeg feature extractor and feature on the target image, indicates the direction and position at new views. In our methodology, we leverage both photometric loss and semantic loss to supervise the generation of rasterized new views. In order for geometry prediction and semantic feature lifting, we employ a confidence-weighted depth loss applied to the input views. The parameters , , are set to 0.25, 0.3, and 1.5, respectively, as determined by the grid search.
@@ 4 Experiments
| NeRF-DFF | 20.52s | 1min2s | 0.4540 | 0.7173 | 27.6806 | 9.6159 | 0.4037 | 0.6755 | 19.8681 | 0.6650 | 0.3629 |
| Feature-3DGS | 20.52s | 18mins36s | 0.4453 | 0.7276 | 12.9595 | 21.0732 | 0.4223 | 0.7174 | 24.4998 | 0.8132 | 0.2293 |
| pixelSplat | 20.52s | 0.064s | - | - | - | - | - | - | 24.8922 | 0.8392 | 0.1641 |
| Ours | 0.108s | 0.5034 | 0.7740 | 3.3853 | 67.7789 | 0.5078 | 0.7686 | 24.3996 | 0.8072 | 0.2506 | |
@@ 4.1 Implementation Details
For our architecture, we employ ViT-Large as the encoder and ViT-Base as the decoder, complemented by a DPT head [42] for pixel-wise geometry regression. We initialize the geometry prediction layers using DUSR3r [19]. Point Transformer layers consists of 5 encoder and 4 decoder blockers with progressive downsampling and upsamping. The cross-model fusion strategy is implemented at the output of the last encoder and the output of the first decoder. The entire system is optimized end-to-end using the loss function described in Eq. 5. The training of our model contains 100 epochs, leveraging a combined dataset of ScanNet++[48] and Scannet[49], of 1565 scenes. Training is on 8 Nvidia A100 GPU lasts for 3 days. We start with a base learning rate of 1e-4 and incorporate a 10-epoch warm-up period. AdamW is employed as the optimizer for all experiments. Evaluation is conducted on 40 unseen scenes from ScanNet. Additionally, we assess on tasks: novel view synthesis, multi-view depth prediction, and language-based semantic segmentation.
@@ 4.2 Semantic 3D Reconstruction
Novel view synthesis is evaluated using NeRF-DFF [12] and Feature-3DGS [14], both of which are capable of predicting RGB values as well as features. In addition, we compared our approach with the state-of-the-art, generalizable, pose-based 3D Gaussian Splatting method, pixelSplat [17], which generates point-based representations through a feed-forward pass. Unlike our method, these existing approaches rely on known camera intrinsics and poses prior to evaluation. As indicated in Table 1, NeRF-DFF and Feature-3DGS tend to overfit on each individual scene, requiring significantly more time than our method, yet performing comparably in terms of output quality. pixelSplat utilizes an Epipolar Transformer, searching along the epipolar line using GT camera parameters to regress Gaussian attributes, resulting in longer inference times. Visualizations in Figure 4 demonstrate that our results are sharper and exhibit fewer artifacts than NeRF-DFF, handle lighting changes more effectively than pixelSplat, and are comparable to Feature-3DGS in performance.
@@ 4.3 Ablation Studies
| [1] | Baseline | 0.4562 | 0.6940 | 24.0006 | 0.7981 |
| [2] | [1] + Fuse Encoder Feat. | 0.5410 | 0.8083 | 23.6723 | 0.7876 |
| [3] | [2] + Fuse LSeg Feat. | 0.5586 | 0.8505 | 23.8585 | 0.7902 |
| [4] | [3] + Multi-scale Fusion | 0.6042 | 0.8681 | 24.3996 | 0.8072 |
@@
Incorporating the encoder feature from ViT into the hidden layer of the point-aggregation layer (Sec. 3.2) demonstrates that such cross-model information flow significantly benefits the segmentation task, improving the mean Intersection over Union (mIoU) from 0.4562 to 0.5410 (Exp #1 2).
@@
Employing cross-model fusion, where latent features of the semantic model are integrated into the middle layers of point-based aggregation, improves injection of semantically rich embeddings (0.5410 to 0.5586, Exp # 2 3). The decoded features confirm that the lifted feature field produces higher-quality feature maps, with the semantic mIoU improving from 0.5586 to 0.6042 (Exp #3 4) through multi-scale fusion.

### v2

@@ Abstract
In this work, we introduce the Large Spatial Model (LSM), which directly processes unposed RGB images into semantic radiance fields. LSM simultaneously estimates geometry, appearance, and semantics in a single feed-forward pass and can synthesize versatile label maps by interacting through language at novel views. Built on a general Transformer-based framework, LSM predicts global geometry via pixel-aligned point maps. To improve spatial attribute regression, we adopt local context aggregation with multi-scale fusion, enhancing the accuracy of fine local details. To address the scarcity of labeled 3D semantic data and enable natural language-driven scene manipulation, we incorporate a pre-trained 2D language-based segmentation model into a 3D-consistent semantic feature field. An efficient decoder parameterizes a set of semantic anisotropic Gaussians, allowing supervised end-to-end learning. Comprehensive experiments on various tasks demonstrate that LSM unifies multiple 3D vision tasks directly from unposed images, achieving real-time semantic 3D reconstruction for the first time.
@@ 1 Introduction
Furthermore, reasoning about and interacting with the environment would benefit from a comprehensive 3D understanding. Open-vocabulary methods, which perform semantic segmentation without relying on a fixed set of labels, provide notable flexibility. However, unlike single-image understanding, the absence of large-scale and diverse 3D scene data with accurate multiview language annotations complicates the challenge. Efforts have been made to integrate 2D features into frameworks such as NeRF [12, 13, 14] and 3D-GS [15, 16]. Yet, these methods, such as Feature-3DGS [15] , typically require overfitting each 3D scene separately with extensive captured viewpoints and preprocessing camera poses using Structure-from-Motion.
To address the challenges outlined above, we propose for the first time a novel unified framework for these key 3D vision subproblems: dense 3D reconstruction, open-vocabulary semantic segmentation, and novel view synthesis from unposed and uncalibrated images. Our approach leverages a single Transformer-based model that learns the attributes of a 3D scene via semantic anisotropic Gaussians. Unlike previous methods that rely on epipolar Transformers with known camera parameters [17, 18, 19] or require extensive per-scene fitting [5, 15], we employ a coarse-to-fine strategy. This strategy predicts dense 3D geometry using pixel-aligned point maps, progressively refining these points into anisotropic Gaussians in a single feed-forward pass.
@@ 2 Related Work
Structure-from-Motion (SfM) aims to jointly estimate camera poses and reconstruct sparse 3D structures from multiple views. Traditional pipelines [1] involve multiple stages, including descriptor extraction, correspondence estimation, and incremental bundle adjustment. Recent advances in learning-based techniques [21, 22, 23, 24, 25] have further improved the accuracy and efficiency of SfM. These methods are widely adopted in 3D vision tasks, where differentiable neural representations typically assume accurate camera poses provided by SfM. For instance, NeRF [4] and its successors [26] rely on poses estimated offline via COLMAP [1, 27]. Similarly, 3D Gaussian Splatting [5] uses SfM-generated 3D points for initialization and has been applied to robotics [28, 29, 30], healthcare [31, 32, 33], and many other domains [34, 35, 36]. Beyond novel view synthesis, lifting 2D features to 3D has gained traction in various editing tasks [13, 15, 37, 14].
@@
3D reconstruction is a long-standing problem in computer vision, with traditional approaches like SfM [38, 39, 1], Multi-view Stereo (MVS) [3, 2, 40, 41], and Signed Distance Function (SDF) [42, 43]. More recent techniques utilize neural representations, including implicit [4] and explicit [5] formats to generate 3D models. Semantic understanding is often integrated during the reconstruction process [44], or through additional optimization steps [12, 13]. However, most methods depend on a preprocessing step like SfM [1] to estimate camera calibration, poses, and sparse point clouds before dense reconstruction, either through feed-forward prediction or test-time optimization. This reliance on calibration and pose estimation limits scalability with large-scale data, contrasting the success seen with large foundation models [45]. The latest pose-free, feedforward approaches, such as Scene Representation Transformers[46, 47, 48], have advanced the concept of representing multiple images as a “set latent scene representation,” allowing for novel view generation even in the presence of inaccurate camera poses or without any pose information. However, these methods struggle to produce explicit geometry. DUSt3R[20] addresses this limitation by predicting dense point clouds directly from unposed stereo image pairs, enabling pixel-aligned geometry prediction at normalized scales. Practically, dense point prediction requires accurate multi-view RGB-D pairs, which significantly limits its scalability. InstantSplat[49] addresses this by utilizing novel-view synthesis with only posed image data and employing Gaussian Bundle Adjustment to jointly optimize camera and scene parameters for ultra-fast dense 3D reconstruction.
In contrast, our framework offers a holistic solution for dense 3D semantic reconstruction from unposed images. It integrates dense 3D geometry reconstruction, and language-based 3D interaction, while minimizing the need for extensive data annotation by using novel view synthesis as a core task. Since dense 3D annotations are often scarce in real-world scenarios, we propose semantic anisotropic Gaussians to lift 2D features map to 3D semantic embeddings without additional annotation. Our approach addresses higher-level 3D tasks in perception and dense 3D reconstruction compared to DUSt3R [20], by utilizing lightweight annotations and solving these tasks jointly within a unified framework.
@@ 3.1 Dense Geometry Prediction
where is pixel-aligned confidence map, same as DUSt3R, indicates all valid points to the origin, means view index relative to view index 1, is a hyper-parameter that apply regularization, encouraging the network to perform robustly in challenging areas. The depth error is calculated by
where the normalization factors (z and ) indicate that the predicted and ground-truth pointmaps are processed by normalizing as follows:
@@ 3.2 Point-wise Feature Aggregation
Building on the foundational work in NeRF[4] and Multi-view Stereo[3], which employ a coarse-to-fine strategy for high-quality radiance field and depth estimation, we also aggregate the initial predicted geometry by applying a Transformer [54] at the point level, leveraging hierarchical representations to achieve more refined, point-based regression.
@@
Rather than relying solely on a single network to represent the scene, we employ two Transformer-based networks optimized for distinct tasks: one for capturing “coarse” global geometry and another for “fine” local information aggregation. Initially, we integrate stereo point maps, including color information for each point primitive, formulated as to serve as input. Unlike tokenized image patches, point primitives carry distinct geometric significance within Euclidean space. Inspired by recent advancements in point-cloud processing [55, 56, 57], we employ a Transformer within a localized window to perform point-wise aggregation, selectively emphasizing key features from neighboring primitives. Point-wise encoding and decoding are essential for refining scene representation, utilizing multiscale aggregation across five hierarchical levels.
@@
The original point features contain explicit and precise spatial information, which is critical for accurate geometry reconstruction. In contrast, the image token features , from image encoder (Sec. 3.1) are rich in semantic content, providing important contextual information that enhances general understanding of the scene. Cross-model fusion enables the integration of detailed spatial geometry with semantic richness:
@@ 3.3 Learning Hierarchical Semantics
To facilitate semantic 3D representation, we augment the anisotropic 3D Gaussians with a learnable semantic feature embedding (a.k.a. semantic anisotropic Gaussians) and rasterize into the 2D image plane by blending Gaussians that overlap with each pixel using a feature rendering function.
indicates the final rasterized feature embedding on image plane, and is semantic embedding on anisotropic Gaussians.
@@
After obtaining , we optimize by minimizing the difference between the rasterized feature map and the feature maps generated by a pre-trained 2D model. Unlike the previous method [15] which requires test time optimization, we transform the estimation of the feature field into a fully learnable process.
Feature maps () from a pre-trained 2D multi-modal model [58] are inherently view-inconsistent due to the lack of spatial awareness during the model’s training. To elevate multi-view feature embeddings into a coherent 3D feature field for holistic 3D understanding, we introduce a dynamic fusion strategy employing an attention-based correlation module. This module is specifically designed to learn blending weights for each token within Point Transformer [54] from the input pixel-wise feature embedding (). We employ attention blocks as described in Eq.3.2 to synchronize in the latent spaces through a supplementary set of cross-attention layers. The visual feature from LSeg[58], denoted as , is utilized for this purpose. This loss function is minimized during training by utilizing rasterized feature maps on new views and directly inferred feature maps using ground truth images on new views (LSeg [58]), thereby facilitating the learning of blending weights for consistent semantic field regression.
@@
To improve model efficiency, we propagate information from ViT Encoder feature and the frozen semantic feature , to the 3D latent space (point feature ) which has fewer tokens, thereby enabling selective attention to critical features. We further refine feature fusion across multiple stages, optimizing information flow while minimizing additional computational overhead. Novel view synthesis serves as an effective task to encode the complete geometric and appearance features into a low-dimensional 3D latent space, while recovering a set of semantic anisotropic Gaussians through learning from large-scale data and end-to-end training.
@@ 3.4 Training Objective
where and are rasterized and GT pixel intensities, denotes represented 3D scene using a set of 3D semantic anisotropic Gaussians, and denotes rendered LSeg feature extractor and feature on the target image, indicates the direction and position at new views. In our methodology, we leverage both photometric loss and semantic loss to supervise the generation of rasterized new views. In order for geometry prediction and semantic feature lifting, we employ a confidence-weighted depth loss applied to the input views. The parameters , , are set to 0.25, 0.3, and 1.5, respectively, as determined by the grid search.
@@ 4 Experiments
NeRF-DFF | 20.52s | 1min2s | 0.4540 | 0.7173 | 27.68 | 9.61 | 0.4037 | 0.6755 | 19.86 | 0.6650 | 0.3629 |
Feature-3DGS | 20.52s | 18mins36s | 0.4453 | 0.7276 | 12.95 | 21.07 | 0.4223 | 0.7174 | 24.49 | 0.8132 | 0.2293 |
pixelSplat | 20.52s | 0.064s | - | - | - | - | - | - | 24.89 | 0.8392 | 0.1641 |
Ours | 0.108s | 0.5034 | 0.7740 | 3.38 | 67.77 | 0.5078 | 0.7686 | 24.39 | 0.8072 | 0.2506 |
@@ 4.1 Implementation Details
For our architecture, we employ ViT-Large as the encoder and ViT-Base as the decoder, complemented by a DPT head [53] for pixel-wise geometry regression. We initialize the geometry prediction layers using DUSt3R [20]. Point Transformer layers consists of 5 encoder and 4 decoder blocks with progressive downsampling and upsamping. The cross-model fusion strategy is implemented at the output of the last encoder and the output of the first decoder. The entire system is optimized end-to-end using the loss function described in Eq. 6. The training of our model contains 100 epochs, leveraging a combined dataset of ScanNet++[60] and Scannet[61], of 1565 scenes. Training is on 8 Nvidia A100 GPU lasts for 3 days. We start with a base learning rate of 1e-4 and incorporate a 10-epoch warm-up period. AdamW is employed as the optimizer for all experiments. Evaluation is conducted on 40 unseen scenes from ScanNet. Additionally, we assess on tasks: novel view synthesis, multi-view depth prediction, and 3D language-based semantic segmentation.
@@ 4.2 Semantic 3D Reconstruction
Novel view synthesis is evaluated using NeRF-DFF [13] and Feature-3DGS [15], both of which are capable of predicting RGB values as well as features. In addition, we compared our approach with the state-of-the-art, generalizable, pose-based 3D Gaussian Splatting method, pixelSplat [18], which generates point-based representations through a feed-forward pass. Unlike our method, these existing approaches rely on known camera intrinsics and poses prior to evaluation. As indicated in Table 1, NeRF-DFF and Feature-3DGS tend to overfit on each individual scene, requiring significantly more time than our method, yet performing comparably in terms of output quality. pixelSplat utilizes an Epipolar Transformer, searching along the epipolar line using GT camera parameters to regress Gaussian attributes, resulting in longer inference times. Visualizations in Figure 4 demonstrate that our results are sharper and exhibit fewer artifacts than NeRF-DFF, and are comparable to Feature-3DGS and pixelSplat in performance.
@@ 4.3 Ablation Studies
[1] | Baseline | 0.4562 | 0.6940 | 24.00 | 0.7981 |
[2] | [1] + Fuse Encoder Feat. | 0.5410 | 0.8083 | 23.67 | 0.7876 |
[3] | [1] + Fuse LSeg Feat. | 0.5586 | 0.8505 | 23.85 | 0.7902 |
[4] | [1] + [2] + [3] + Multi-scale Fusion | 0.6042 | 0.8681 | 24.39 | 0.8072 |
@@
Incorporating the geometry encoder feature from ViT into the hidden layer of the point-aggregation layer (Sec. 3.2) demonstrates that such cross-model information flow significantly benefits the segmentation task, improving the mean Intersection over Union (mIoU) from 0.4562 to 0.5410 (Exp #1 2).
@@
Employing cross-model fusion, where latent features of the semantic model are integrated into the middle layers of point-based aggregation, also improves injection of semantically rich embeddings (0.4562 to 0.5586, Exp # 1 3). The decoded features confirm that the lifted feature field produces higher-quality feature maps, with the semantic mIoU improving from 0.5586 to 0.6042 (Exp #3 4) through multi-scale fusion.
Module Timing.
We analyze the computational cost of each module by running inference 1,000 times on the ScanNet test dataset with the model, as shown in Table 3, and calculating the average inference time for each module of the Large Spatial Model.
Module | Inference Time (seconds) |
Dense Geometry Prediction (Sec. 3.1) | 0.029 |
Point-wise Aggregation (Sec. 3.2) | 0.046 |
Feature Lifting (Sec. 3.3) | 0.019 |
Total | 0.096 |
Evaluation of Generalizable Methods on New Datasets.
To avoid potential overfitting, we adopt the Replica dataset[63], a photorealistic simulated 3D dataset with accurate RGB, dense depth maps, and semantic annotations for comprehensive evaluation. We use the same data preparation with Feature-3DGS. LSM generalizes well to the simulated Replica test set, achieving the best depth estimation metrics and enabling 3D semantic segmentation, which is unique among generalizable methods. Splatter Image[64], an ultra-fast monocular 3D object reconstruction method using 3D-GS, performs well for object reconstruction with masked backgrounds but struggles with scene-wise reconstruction in complex backgrounds.
Model | Offline SfM | mIoU | rel | PSNR | |
Replica | pixelSplat | Required | None | 20.14 | 26.28 |
Splatter Image | Required | None | None | 12.37 | |
Ours | Not Required | 0.51 | 4.91 | 23.10 |